• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 04:04:13 on October 7,2010

    Entry number 333845

    keyword=Compton tune not successful

    After many hours, Yves has given up Compton tuning for tonight. He says the tune in the hall A line is as good as possible, so hopefully the new injector tune (for running with all 3 halls) will fix our Compton rates.

    Right now, with the HV=-1450, and the detector at x=70.3, y=0 (out of the Compton beamline), the rates are 430Hz/uA (at 5uA) ... at this HV and out of the beamline we should see no rate at all.


    FIGURE 1